In the coordination compound, K4[Ni(CN)4], the oxidation state of nickel is

Grade:12

3 Answers

Gaurav
9 years ago
Hello Student
In the coordination compound, K4[Ni(CN)4], the oxidation state of nickel is 0.
[Ni(CN)4]4-
x + 4(-1) = -4
x = 0
Ravi Kumar
6 years ago
The oxidation number of Ni in complex compound K4[Ni(CN)4] is equal to zero. Explanation ; Let the oxidation number of Ni is `x`Now. K4[Ni(CN)4] =1.4 + X.1 +(-1).4 =0=4 + x - 4 =0=x =0. Proved. , here K=+1,CN = - 1
Naturally Insane
2 years ago

Let oxidation state of Nickel be

(1×4)+x+(4×1)=0
4+x4=0
 x=0
Hence, in the given compound, the oxidation state of Nickel is zero.